OpenStudy (anonymous):

if 8 men can reap 80 hectares in 24 days, how many hectares can 36 men reap in 30 days. please explain step by step

OpenStudy (anonymous):

@satellite73 , @Hero

OpenStudy (whpalmer4):

You want to find the amount that 1 man does in 1 day. 8 men do 80 hectares in 24 days, so 1 man does 1/8 th of that = 10 hectares in 24 days = 10/24 hectare/day. Now multiply that rate by 36 men and 30 days.

OpenStudy (anonymous):

find the unit rate

OpenStudy (anonymous):

unit rate ??

OpenStudy (whpalmer4):

Alternatively, just multiply all the ratios: You know what 8 men do in 24 days, and you have 36/8 men and 30/24 days. Same idea.

OpenStudy (whpalmer4):

Unit rate: amount 1 worker does each day.

OpenStudy (shubhamsrg):

No of men increase from 8 to 36, hence manpower increases by 36/8 No of days from 24 to 30, so now the time taken would be 30/24 Hence no. of hectares will be increased by (36/8)(30/24) or => 80 *(36/8) *(30/24) just simplify

OpenStudy (anonymous):

no off hectors given in the answer is 450

OpenStudy (whpalmer4):

80 * 36 / 8 * 30 / 24 = 10 * 36 * 30 / 24 = 300 * 3 / 2 = 450

OpenStudy (whpalmer4):

Of if you found the unit rate of 10/24 hectares/day, 36 * 30 * 10 / 24 = 450.

OpenStudy (shubhamsrg):

Or, alternative way of looking can be 8 men take 24 days for 80 hectares so 1 man takes 24 days for 10 hecatres so 1 man takes 1 day for 10/24 hecatres so 36 men take 1 day for (10/24)*36 hectares so 36 men take 30 days for (10/24)*36*30 hectares And hence is the ans.

OpenStudy (anonymous):

basically i have test for job and i dont know this type mathematics too much. so can you me the easiest way to solve this type of problem @shubhamsrg

OpenStudy (anonymous):

@whpalmer4

OpenStudy (shubhamsrg):

Just as every one else said here, try to convert into single units,you should always make out if we reduce no. of men or days, will workload inc, dec, likewise

OpenStudy (whpalmer4):

Yeah, when in doubt, find out how much work 1 unit can do (whether that be workers, machines, or whatever), then you can find the answer to how long it will take if you have more time and more workers. The approach we showed with proportions is a bit faster, but probably isn't as safe if you don't really understand what it represents.